OpenStudy (laronjaxon):

Let f(x) = 5x – 26 and g(x) = 6x. Part 1 What is (f ◦ g)(x) ? Part 2 Use complete sentences to describe the method you used to solve this problem.

OpenStudy (anonymous):

f(g(x)) = f(6x) = 5(6x) - 26 = 30x -26